In the Triwizard Tournament hosted by Hogwarts School, Harry Potter walks 1.5 m southeast, 2.5 m 60° west of south and finally 3.5 m straight south to locate the Triwizard Cup hidden in a massive hedge maze. Determine Potter’s final position relative to his starting point. 


1
Expert's answer
2022-02-14T14:38:30-0500

We know that

"OQ=2.5m"

"QT=OQsin\theta"

"QT=2.5sin60=2.16m"

"TP=1.5sin\theta"

"TP=1.5sin45=1.06m"

We know that


"RP=\sqrt{RP^2+QP^2}=\sqrt{3.5^2+3.22^2}=\sqrt{22.62}=4.75m"